# Towards Federated Graph Learning in One-shot Communication

### v6

@@ A.3.4 Performance under Model Heterogeneity
| O-pFGL (Ours) | 47.89 0.11 | 13.11 0.34 | 89.59 0.10 | 45.31 0.36 | 91.16 0.04 | 48.56 0.31 |
A.3.5 Inductive Performance
We conduct experiments on three inductive graph datasets: Flickr, Reddit, and Reddit2, and vary the number of clients. Since Flickr has lower homophily, we use the 2-layer SIGN [13] for better graph learning. With 10 clients, experimental results under Louvain and Metis partitions are shown in Table 2 and Table 13, respectively. With 20 clients, experimental results under Louvain and Metis partitions are shown in Table 12 and Table 14, respectively. Our method is scalable and generally performs best.
@@ A.3.5 Inductive Performance
| O-pFGL (Ours) | 47.08 0.27 | 13.37 0.24 | 88.28 0.01 | 31.01 0.47 | 90.25 0.07 | 30.74 0.25 |
A.3.6 More Datasets
We further provide more experimental results on Computers, Photo, CS, and Physics datasets. Computers and Photo are Amazon co-purchase networks and CS and Physics are Co-author networks. The results under Louvain and Metis partitions are shown in Figure 15 and Figure 16, respectively. Our method consistently outperforms others in terms of both accuracy and F1-macro, which demonstrates the general advantages of our method.
@@ A.3.6 More Datasets
| O-pFGL (Ours) | 88.59 0.13 | 33.01 3.66 | 91.79 0.13 | 42.39 0.81 | 89.96 0.20 | 50.33 1.01 | 94.51 0.09 | 66.41 1.24 |
A.3.7 Different GNN Architectures
To demonstrate the generality of our proposed method under different graph models, we conduct experiments with 3 other prevalent graph models, including 1) GraphSage [19], 2) GAT [45], and 3) SGC [49]. The experimental results under Louvain and Metis partitions are shown in Table 17 and Table 18, respectively. Our method shows consistent advantages over other methods.
@@ A.3.7 Different GNN Architectures
| O-pFGL (Ours) | 78.48 | 44.59 | 69.55 | 48.17 | 69.20 | 36.57 | 80.68 | 50.02 | 71.91 | 49.03 | 69.98 | 37.83 | 79.96 | 44.48 | 71.58 | 46.87 | 70.26 | 40.10 |
A.3.8 Hyper-parameters Analysis
@@
Impact of the pre-set number of nodes in synthesized global pseudo graph.
When increasing the number, the generated global pseudo-graph could have more potential to capture more information. But a large number of nodes would incur optimization difficulties and high communication costs when distributing them to clients. Thus, our principles to choose the number are (1) set the smallest number to capture enough information, (2) keep a similar label distribution with the training set. According to Section A.2, we pre-set the number of nodes to synthesize to 1 for three small graphs and pre-set the number of nodes to synthesize to percent of the total amount of nodes in corresponding class in the training set, while we keep as small as possible. We conduct experiments on 10 clients under the Louvain partition to show how the number of nodes impacts the performance. The experimental results are shown in Table 19 and Table 20. For small-scale graphs: Cora, CiteSeer, and PubMed, we pre-set the smallest number of nodes for each class (i.e., one node for each class) and find that it’s already capable. Increasing the number of nodes would not bring benefit since it incurs optimization difficulty. For larger graph: ogbn-arxiv, we pre-set the number of nodes for each class as 0.25% of the total number of corresponding class’s nodes in the train set. Fewer nodes (0.1%, 0.05%) could not fully capture the information in the aggregated class-wise statistics, and lead to a performance drop. More nodes (0.5%, 1%) are more capable but incur more computation and communication costs, and the performance improvement is marginal. Note that the number of nodes in the global pseudo-graph is always small. Even on the largest graph ogbn-product, the total number of nodes of all classes is less than 150. Thus, the memory and communication costs are negligible.
@@
A.3.9 Different Adjacency Matrix Implementations
In our method, we derive the adjacency matrix of the generated global pseudo graph by Eq. 10. We implement the as a simple 3-layer MLP with input dimensions are ( is the node feature dimension) and output dimension is 1 (output , represents the probability of existence of a link). Here we compare the other two different implementations of : (1) make a learnable matrix and directly optimize , (2) is determined by the dot product of (normalized) node features. The experimental results are shown in Table 23. The advantages of the other two designs cannot be generalized to most datasets. Our design, in contrast, offers flexibility via the learnable MLP and value, thus show overall performance advantages. Note that when setting , our design fall back to derive a global pseudo graph where all nodes are isolated.
| Cora | CiteSeer | PubMed | ogbn-arxiv | |||||
| Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | |
| total # nodes = 339 | class1 (# nodes = 50) | class2 (# nodes = 14) | ||||
| F1-macro | Recall | Precision | F1-macro | Recall | Precision | |
| Standalone | 51.28 | 50.00 | 52.63 | 0 | 0 | 0 |
| FedAvg | 45.71 | 40.00 | 53.33 | 0 | 0 | 0 |
| FedPUB | 55.81 | 60.00 | 52.17 | 0 | 0 | 0 |
| FedGTA | 50.00 | 55.00 | 45.83 | 29.63 | 57.14 | 20.00 |
| FedTAD | 42.42 | 35.00 | 53.85 | 0 | 0 | 0 |
| FedSpray | 50.00 | 56.25 | 45.00 | 0 | 0 | 0 |
| FedLoG | 47.06 | 40.00 | 57.14 | 0 | 0 | 0 |
| DENSE | 48.65 | 45.00 | 52.94 | 0 | 0 | 0 |
| Co-Boost | 42.42 | 35.00 | 53.85 | 0 | 0 | 0 |
| O-pFGL (Ours) | 63.83 | 75.00 | 55.56 | 50.00 | 42.86 | 60.00 |
A.3.10 Fine-grained Experimental Results
A.4 Complexity Analysis
@@
We provide the communication analysis of the methods in Table 25. FedAvg, FedPUB, FedTAD, DENSE, and Co-Boost all need to upload and download the whole model parameters (, where is the hidden dimension). FedGTA needs to upload additional mixed moments of neighboring features (, where is the order of the moment and is the propagation steps of Label Propagation) for personalized aggregation. FedSpray needs to upload and download the global feature-structure encoder and structure proxy (, where is the output dimension of the encoder and dimension of the structure proxy). FedLoG needs to upload the pre-trained prompt generator (for simplicity, we use to denote the size of the prompt generator), head-branch classifier and tail-branch classifier (we use and to denote the size of the two branch classifiers), synthetic graph (, where is the pre-set number of nodes), along with the model, to the server. Each client in FedLoG needs to download the aggregated head-branch and tail-branch classifier and the model. The PG, Clfh, and Clft are implemented as multi-layer perceptrons with relatively large hidden dimensions. Thus, FedLoG would incur high upload and download communication costs. In contrast, our method only needs to upload the class-wise statistics () and download the generated global pseudo-graph (). is the number of classes. is the propagation step in Eq. 5 and in our experiments. Although in some rare cases where is extremely large, the upload costs still would not become a huge burden compared with the model size. is the number of nodes in the global pseudo-graph, which is usually a pre-set small value, as indicated in Sec A.2. The communication costs of our method are independent of the model size. The larger the model, the more communication-efficient our method is compared to other methods.
@@
| O-pFGL (Ours) |
Computation Analysis.
We provide computation analysis of methods in Table 26. We slightly abuse the notation by redefining the number of clients as , the number of edges in local graph data as , and the number of nodes in local graph data as . is the number of proxy graphs generated in FedPUB, and is the number of nodes in each proxy graph. is the diffusion step in FedTAD, and is the complexity to calculate the topology embedding, which could be highly computationally expensive. For FedTAD, DENSE, and Co-Boost, is the complexity of constructing the generated graph. is the number of layers of the generator, is the dimension of the latent features in the generator, and is the number of generated nodes. is the complexity of forwarding the generator in the server. For FedLoG, the additional computation costs come from the prompt generator pretraining and graph synthesis. Although the prompt generator and classifiers are multilayer perceptrons, the pre-training process involves bilevel optimization, which would incur high computation costs in practice. For our method, is the complexity of aggregating the statistics. is the estimated number of edges in the generated global pseudo-graph, usually a small value. is the complexity of calculating alignment loss. is the complexity of calculating the smoothness loss. The computation complexity on clients in our method is kept as least as possible, the same as the vanilla FedAvg and Standalone.
@@
We provide memory analysis of methods in Table 27. Standalone, FedAvg, FedPUB, DENSE, and Co-Boost only need to store the model parameters (). FedTAD needs to additionally store the topology embedding calculated by graph diffusion (), thus it faces OOM on larger graphs. FedSpray needs to store additional encoders and structure proxy. FedLoG needs to store the additional prompt generator, two branch classifiers, and generated graphs. Notably, in prompt generator pretraining, FedLoG needs additionally obtain the k-hop subgraph of each node in the local graph, which would incur huge memory (, where is the average degree and is the pre-set number of hops). This is the key factor leading to the out-of-memory issue in dense graphs or large-scale graphs. For FedTAD, DENSE, Co-Boost, and our method, the server needs to additionally store the generated graph, which costs approximately memory. Note that is always a small number. Even on the largest graph ogbn-product, the total number of nodes of all classes is less than 150 (). Thus, it would not incur much memory overhead on the server.
A.5 An Equivalent Server-Efficient Variant Method
@@
An equivalent variant of our method is to move the global pseudo-graph generation process to the clients. Specifically, the server only securely aggregates the uploaded statistics and then distributes to clients (with Homomorphic Encryption [1]). Then the clients generate the global pseudo-graph with aggregated statistics locally. It still supports model heterogeneity and Secure Aggregation protocols. This variant enables extreme efficiency on the server since the server only needs to perform a series of weighted average operations, moving the computation overhead of global pseudo-graph generation to the client-side.
A.6 Related Works
A.6.1 Federated Graph Learning
@@
With the rapid development of federated learning methods, recent works introduce federated graph learning to collaboratively train graph models [31] in a privacy-preserving manner and apply it to many applications [48, 61, 53, 44, 7]. From the graph level, each client possesses multiple completely disjoint graphs (e.g., molecular graphs). Recent works [52, 43] mainly focus on the intrinsic heterogeneity among graphs from different clients. From the subgraph level, the graph possessed by each client can be regarded as a part of a larger global graph. To cope with heterogeneous graph data, GraphFL [46] adopts meta-learning [12] for better generalization. FedGL [5] uploads the node embedding and prediction for global supervision but faces a heavy communication burden and potential privacy concerns. FGSSL [24] augments the local graph to mitigate the heterogeneity. To enhance the model utility on each client, personalized federated graph learning methods are proposed. FedPUB [2] generates random graphs [22] to measure the similarity in model aggregation and conducts adaptive weight masks for better personalization. FedGTA [33] proposes topology-aware personalized optimization. AdaFGL [32] studies the structure non-IID problem. FedTAD [69] utilized ensemble local models to perform data-free distillation on the server. To complete missing connections between graphs of clients, FedSage [63] and FedDEP [64] additionally train a neighborhood generator. FedGCN [57] additionally uploads and downloads the encrypted neighbor features to supplement the features of the neighborhood. To better represent the minority in local graph data, FedSpray [15] learns local class-wise structure proxies to mitigate biased neighboring information. But it needs numerous communication rounds for optimization. FedLoG [30] synthesizes node features in each client, and the server aggregates these node features to generate global synthetic data. Each client then generalizes its local training via the synthetic graph.
A.6.2 One-shot Federated Learning
@@
One-shot federated learning largely reduces communication costs and circumvents potential man-in-the-middle attacks. Mainstream OFL methods can be classified into 3 categories. (1) Ensemble-based: The original OFL study [18] ensembles local models and conducts knowledge distillation with public data. DENSE [62] employs model inversion [58] from the ensemble model to generate images for distillation. FedOV [10] introduces placeholders in the model prediction layer. IntactOFL [60] trains a MoE [27] network by the generated images. Co-Boost [9] further optimizes the generated images and ensemble weights iteratively. (2) Distillation-based: DOSFL [66] and FedD3 [42] conduct dataset distillation locally and upload distilled data for server-side training. (3) Generative-based: FedCVAE [21] trains VAEs for each client to generate similar images, addressing data heterogeneity. FedDISC [54] and FedDEO [55] leverage the pre-trained Stable Diffusion [38] to generate images and mitigate data heterogeneity. However, existing OFL methods primarily focus on image data and are either incompatible or ineffective for graph learning.
A.6.3 Critical Analysis of Current Related Work
Existing FGL methods require numerous communication rounds.

### v7

@@ D.5 Performance under Model Heterogeneity
| O-pFGL | 47.890.11 | 13.110.34 | 89.590.10 | 45.310.36 | 91.160.04 | 48.560.31 |
D.6 Inductive Performance
We conduct experiments on three inductive graph datasets: Flickr, Reddit, and Reddit2, and vary the number of clients. Since Flickr has lower homophily, we use the 2-layer SIGN (Frasca et al. 2020) for better graph learning. With 10 clients, experimental results under Louvain and Metis partitions are shown in Table 13 and Table 15, respectively. With 20 clients, experimental results under Louvain and Metis partitions are shown in Table 14 and Table 16, respectively. Our method is scalable and generally performs best.
@@ D.6 Inductive Performance
| O-pFGL | 47.080.27 | 13.370.24 | 88.280.01 | 31.010.47 | 90.250.07 | 30.740.25 |
D.7 More Datasets
We further provide more experimental results on Computers, Photo, CS, and Physics datasets. Computers and Photo are Amazon co-purchase networks, and CS and Physics are Co-author networks. The results under Louvain and Metis partitions are shown in Figure 17 and Figure 18, respectively. Our method consistently outperforms others in terms of both accuracy and F1-macro, which demonstrates the general advantages of our method.
@@ D.7 More Datasets
| O-pFGL | 88.590.13 | 33.013.66 | 91.790.13 | 42.390.81 | 89.960.20 | 50.331.01 | 94.510.09 | 66.411.24 |
D.8 Different GNN Architectures
To demonstrate the generality of our proposed method under different graph models, we conduct experiments with 3 other prevalent graph models, including 1) GraphSage (Hamilton, Ying, and Leskovec 2017), 2) GAT (Veličković et al. 2017), and 3) SGC (Wu et al. 2019). The experimental results under Louvain and Metis partitions are shown in Table 19 and Table 20, respectively. Our method shows consistent advantages over other methods.
@@ D.8 Different GNN Architectures
| O-pFGL | 78.48 | 44.59 | 69.55 | 48.17 | 69.20 | 36.57 | 80.68 | 50.02 | 71.91 | 49.03 | 69.98 | 37.83 | 79.96 | 44.48 | 71.58 | 46.87 | 70.26 | 40.10 |
| total # nodes = 339 | class1 (# nodes = 50) | class2 (# nodes = 14) | ||||
| F1-macro | Recall | Precision | F1-macro | Recall | Precision | |
| Standalone | 51.28 | 50.00 | 52.63 | 0 | 0 | 0 |
| FedAvg | 45.71 | 40.00 | 53.33 | 0 | 0 | 0 |
| FedPUB | 55.81 | 60.00 | 52.17 | 0 | 0 | 0 |
| FedGTA | 50.00 | 55.00 | 45.83 | 29.63 | 57.14 | 20.00 |
| FedTAD | 42.42 | 35.00 | 53.85 | 0 | 0 | 0 |
| FedSpray | 50.00 | 56.25 | 45.00 | 0 | 0 | 0 |
| FedLoG | 47.06 | 40.00 | 57.14 | 0 | 0 | 0 |
| FedGM | 56.42 | 55.00 | 57.89 | 0 | 0 | 0 |
| GHOST | 48.65 | 45.00 | 52.94 | 0 | 0 | 0 |
| DENSE | 48.65 | 45.00 | 52.94 | 0 | 0 | 0 |
| Co-Boost | 42.42 | 35.00 | 53.85 | 0 | 0 | 0 |
| O-pFGL | 63.83 | 75.00 | 55.56 | 50.00 | 42.86 | 60.00 |
D.9 Fine-grained Experimental Results
Appendix E Hyper-parameters Analysis
@@
Impact of the pre-set number of nodes in synthesized global surrogate graph.
When increasing the number, the generated global surrogate graph could have more potential to capture more information. But a large number of nodes would incur optimization difficulties and high communication costs when distributing them to clients. Thus, our principles to choose the number are (1) set the smallest number to capture enough information, (2) keep a similar label distribution with the training set. According to Section C, we pre-set the number of nodes to synthesize to 1 for three small graphs and pre-set the number of nodes to synthesize to percent of the total number of nodes in the corresponding class in the training set, while we keep as small as possible. We conduct experiments on 10 clients under the Louvain partition to show how the number of nodes impacts the performance. The experimental results are shown in Table 22 and Table 23. For small-scale graphs: Cora, CiteSeer, and PubMed, we pre-set the smallest number of nodes for each class (i.e., one node for each class) and find that it’s already capable. Increasing the number of nodes would not bring benefit since it incurs optimization difficulty. For larger graph: ogbn-arxiv, we pre-set the number of nodes for each class as 0.25% of the total number of corresponding class’s nodes in the train set. Fewer nodes (0.1%, 0.05%) could not fully capture the information in the aggregated class-wise statistics, and lead to a performance drop. More nodes (0.5%, 1%) are more capable but incur more computation and communication costs, and the performance improvement is marginal. Note that the number of nodes in the global surrogate graph is always small. Even on the largest graph ogbn-product, the total number of nodes of all classes is less than 150. Thus, the memory and communication costs are negligible.
@@
Appendix F Complexity Analysis
@@
We provide the communication analysis of the methods in Table 26. FedAvg, FedPUB, FedTAD, DENSE, and Co-Boost all need to upload and download the whole model parameters (, where is the hidden dimension). FedGTA needs to upload additional mixed moments of neighboring features (, where is the order of the moment and is the propagation steps of Label Propagation) for personalized aggregation. FedSpray needs to upload and download the global feature-structure encoder and structure proxy (, where is the output dimension of the encoder and dimension of the structure proxy). FedLoG needs to upload the pre-trained prompt generator (for simplicity, we use to denote the size of the prompt generator), head-branch classifier and tail-branch classifier (we use and to denote the size of the two branch classifiers), synthetic graph (, where is the pre-set number of nodes), along with the model, to the server. Each client in FedLoG needs to download the aggregated head-branch and tail-branch classifier and the model. The PG, Clfh, and Clft are implemented as multi-layer perceptrons with relatively large hidden dimensions. Thus, FedLoG would incur high upload and download communication costs. FedGM needs to upload the synthesized graphs and a link predictor (PGE) on each client. The link predictor is usually larger than the model. And the server in FedGM would distribute the aggregated synthesized graphs to clients. For GHOST, each client needs to upload the proxy model, which is inherently a generator G. And the server would distribute the generalized model. In contrast, our method only requires uploading the class-wise statistics () and downloading the generated global surrogate graph (). is the number of classes. is the propagation step in Eq. 4 and in our experiments. Although in some rare cases where is extremely large, the upload costs would still not become a huge burden compared to the model size. is the number of nodes in the global surrogate graph, which is usually a pre-set small value, as indicated in Sec C. The communication costs of our method are independent of the model size. The larger the model, the more communication-efficient our method is compared to other methods.
@@
| O-pFGL |
Computational Analysis.
We provide computational analysis of methods in Table 27. We slightly abuse the notation by redefining the number of clients as , the number of edges in local graph data as , and the number of nodes in local graph data as . is the number of proxy graphs generated in FedPUB, and is the number of nodes in each proxy graph. is the diffusion step in FedTAD, and is the complexity to calculate the topology embedding, which could be highly computationally expensive. For FedTAD, DENSE, and Co-Boost, is the complexity of constructing the generated graph. is the number of layers of the generator, is the dimension of the latent features in the generator, and is the number of generated nodes. is the complexity of forwarding the generator in the server. For FedLoG, the additional computation costs come from the prompt generator pretraining and graph synthesis. Although the prompt generator and classifiers are multilayer perceptrons, the pre-training process involves bilevel optimization, which would incur high computation costs in practice. For FedGM, each client needs to conduct gradient matching to synthesize a graph. The server needs to aggregate and infer the synthesized graphs with uploaded PGE from clients. For GHOST, the computational complexity is very high. Each client needs to construct a pseudo graph with the same size as the local graph by feature matrix multiply. This operation is , which would be impractical on large-scale datasets. On the server side, to integrate the knowledge and avoid catastrophic forgetting, the server needs to traverse every edge in the pseudo graphs. The number of edges is approximately the same order as the number of edges in the latent global graphs. Thus, this process is with complexity, also impractical under large-scale datasets. For our method, is the complexity of aggregating the statistics. is the estimated number of edges in the generated global surrogate graph, usually a small value. is the complexity of calculating alignment loss. is the complexity of calculating the smoothness loss. The computational complexity on clients in our method is kept as least as possible, the same as the vanilla FedAvg and Standalone.
@@
We provide memory analysis of methods in Table 28. Standalone, FedAvg, FedPUB, DENSE, and Co-Boost only need to store the model parameters (). FedTAD needs to additionally store the topology embedding calculated by graph diffusion (), thus it faces OOM on larger graphs. FedSpray needs to store additional encoders and the structure proxy. FedLoG needs to store the additional prompt generator, two branch classifiers, and generated graphs. Notably, in prompt generator pretraining, FedLoG additionally needs to obtain the k-hop subgraph of each node in the local graph, which would incur huge memory (, where is the average degree and is the pre-set number of hops). This is the key factor leading to the out-of-memory issue in dense graphs or large-scale graphs. FedGM needs to store the gradient of synthesized graphs additionally, and the server needs to store the PGE from clients. GHOST needs to store the pseudo graph on both clients and the server, thus incurs high memory consumption with , which is impractical on large-scale datasets. For FedTAD, DENSE, Co-Boost, and our method, the server needs to additionally store the generated graph, which costs approximately memory. Note that is always a small number. Even on the largest graph ogbn-product, the total number of nodes of all classes is less than 150 (). Thus, it would not incur much memory overhead on the server.
Appendix G An Equivalent Server-Efficient Variant
@@
An equivalent variant of our method is to move the global surrogate graph generation process to the clients. Specifically, the server only securely aggregates the uploaded statistics and then distributes to clients (with Homomorphic Encryption (Acar et al. 2018)). Then the clients generate the global surrogate graph with aggregated statistics locally. It still supports model heterogeneity and Secure Aggregation protocols. This variant enables extreme efficiency on the server since the server only needs to perform a series of weighted average operations, moving the computation overhead of global surrogate graph generation to the client side.
Appendix H Related Works
H.1 Federated Graph Learning
@@
With the rapid development of federated learning methods, recent works introduce federated graph learning to collaboratively train graph models (Kipf and Welling 2016) in a privacy-preserving manner and apply it to many applications (Wu et al. 2021; Zhang et al. 2024a; Yan et al. 2024; Tang et al. 2024; Chen et al. 2021b). From the graph level, each client possesses multiple completely disjoint graphs (e.g., molecular graphs). Recent works (Xie et al. 2021; Tan et al. 2023, 2024; Fu et al. 2025) mainly focus on the intrinsic heterogeneity among graphs from different clients. From the subgraph level, the graph possessed by each client can be regarded as a part of a larger global graph. To cope with heterogeneous graph data, GraphFL (Wang et al. 2022) adopts meta-learning (Finn, Abbeel, and Levine 2017) for better generalization. FedGL (Chen et al. 2024) uploads the node embedding and prediction for global supervision but faces a heavy communication burden and potential privacy concerns. FGSSL (Huang et al. 2024) augments the local graph to mitigate the heterogeneity. To enhance the model utility on each client, personalized federated graph learning methods are proposed. FedPUB (Baek et al. 2023) generates random graphs (Holland, Laskey, and Leinhardt 1983) to measure the similarity in model aggregation and conducts adaptive weight masks for better personalization. FedGTA (Li et al. 2024b) proposes topology-aware personalized optimization. AdaFGL (Li et al. 2024a) studies the structure non-IID problem. FedTAD (Zhu et al. 2024b) utilized ensemble local models to perform data-free distillation on the server. FedGM (Zhang et al. 2025) adopts gradient matching to synthesize graphs and uploads the synthesized graphs or gradients to the server for aggregation. GHOST (Qian et al. 2025) trains a proxy model to align with local graph data, and the server integrates the knowledge from the uploaded proxy model. To complete missing connections between graphs of clients, FedSage (Zhang et al. 2021) and FedDEP (Zhang et al. 2024b) additionally train a neighborhood generator. FedGCN (Yao et al. 2024) additionally uploads and downloads the encrypted neighbor features to supplement the features of the neighborhood. FedStruct (Aliakbari, Östman et al. 2024) decouples the structure learning and node representation learning. To better represent the minority in local graph data, FedSpray (Fu et al. 2024) learns local class-wise structure proxies to mitigate biased neighboring information. But it needs numerous communication rounds for optimization. FedLoG (Kim et al. 2025) synthesizes node features in each client, and the server aggregates these node features to generate global synthetic data. Each client then generalizes its local training via the synthetic graph.
H.2 One-shot Federated Learning
@@
One-shot federated learning largely reduces communication costs and circumvents potential man-in-the-middle attacks. Mainstream OFL methods can be classified into 3 categories. (1) Ensemble-based: The original OFL study (Guha, Talwalkar, and Smith 2019) ensembles local models and conducts knowledge distillation with public data. DENSE (Zhang et al. 2022) employs model inversion (Yin et al. 2020) from the ensemble model to generate images for distillation. FedOV (Diao, Li, and He 2023) introduces placeholders in the model prediction layer. IntactOFL (Zeng et al. 2024) trains a MoE (Jacobs et al. 1991) network by the generated images. Co-Boost (Dai et al. 2024) further optimizes the generated images and ensemble weights iteratively. (2) Distillation-based: DOSFL (Zhou et al. 2020) and FedD3 (Song et al. 2023) conduct dataset distillation locally and upload distilled data for server-side training. (3) Generative-based: FedCVAE (Heinbaugh, Luz-Ricca, and Shao 2023) trains VAEs for each client to generate similar images, addressing data heterogeneity. FedDISC (Yang et al. 2024a) and FedDEO (Yang et al. 2024b) leverage the pre-trained Stable Diffusion (Rombach et al. 2022) to generate images and mitigate data heterogeneity. However, existing OFL methods primarily focus on image data and are either incompatible or ineffective for graph learning.
H.3 Critical Analysis of Current Related Work
Most FGL methods are ineffective within one-shot communication under data heterogeneity.